OpenStudy (anonymous):

how would you isolate this? S=u+v/2

OpenStudy (anonymous):

S= u+v over 2

OpenStudy (anonymous):

S=u+(v)/(2)

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

what variable are you trying to isolate?

OpenStudy (anonymous):

oops sorry, for u

OpenStudy (anonymous):

u-v?

OpenStudy (anonymous):

S-(v)/(2)=u

OpenStudy (anonymous):

ohhh...so can you tell me how you did it like the stepds....sorry I have a exam tomorow